OpenStudy (lukecrayonz):

Find the exact values of the five remaining trigonometric functions of theta. tan theta = 2, where sin theta > 0 and cos theta > 0

OpenStudy (anonymous):

theta is located in the first quadrant

OpenStudy (lukecrayonz):

? I have to find cos, sec, sin, cot, and csc, not sure how.

OpenStudy (anonymous):



OpenStudy (lukecrayonz):

Okay well my book is telling me to use the unit circle but that seems a lot easier :O

OpenStudy (anonymous):

\[\sin(\theta) = \frac{opp}{hyp} = \frac{2}{\sqrt{5}} = \frac{2 \sqrt{5}}{5}\]

OpenStudy (anonymous):

\[\cos(\theta) = \frac{adj}{hyp} = \frac{1}{\sqrt{5}} = \frac{\sqrt{5}}{5}\]

OpenStudy (anonymous):

\[\csc(\theta) = \frac{hyp}{opp} = \frac{\sqrt{5}}{2}\]

OpenStudy (anonymous):

\[\sec(\theta) = \frac{hyp}{adj} = \sqrt{5}\]

OpenStudy (lukecrayonz):

i got it now :D

OpenStudy (anonymous):

\[\cot(\theta) = \frac{adj}{hyp} = \frac{1}{2}\]

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

Oops \[\cot(\theta) = \frac{adj}{opp}\] but yeah lol..